OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (anonymous):

Slope is 1/2

OpenStudy (anonymous):

bring the eqution in to y=mx+c form then compare it... hear mis slope...

OpenStudy (anonymous):

it's y=(2x-10)/4 so the slope would be the derivative of this function which is 2/4 = 1/2.